Circular No. 6107

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

1E 0035.4–7230 AND RX J0122.9–7521 P. C. Schmidtke, A. P. Cowley, and T. K. McGrath, Arizona State

University; and J. B. Hutchings and D. Crampton, Dominion Astrophysical Observatory, communicate: “Using data obtained at Cerro Tololo Interamerican Observatory, we report observations of two supersoft x-ray sources in the Small Magellanic Cloud.

1E 0035.4–7230 is found to have an orbital period of 0.1719 day, with minimum light at 1994 Nov. 3.105 UT, based on optical photometry of the star (α = 0^{h}37^{m}19^{s}.8, δ = −72^{o}14′13′′, equinox 2000.0; Orio et al. 1994, A.Ap. 289, L11). The light curve is nearly sinusoidal, with mean V = 20.2 and a range of ∼ 0.4 mag. Average colors are B − V = −0.2 and U − B = −0.8. Spectra show a blue continuum with only weak, variable emission lines of O VI (528.9 nm), occasionally He II (468.6 nm), and a moderate-strength unidentified feature at ∼ 395.0 nm.

RX J0122.9–7521 (Kahabka et al. 1994, A.Ap. 288, 538) has been optically identified with an extremely blue object at α = 1^{h}22^{m}52^{s}.9, δ = −75^{o}21′13′′ (equinox 2000.0), with V = 15.5, B −V = −0.4, U −B = −1.0. The spectrum is peculiar, with only a few emission features, including He II (468.6 nm) and O VI (381.1, 383.4, 528.9 nm). Small changes in line flux and velocity are observed on a timescale of hours or less.”

X-RAY NOVA IN SCORPIUS R. M. Hjellming and M. P. Rupen, National Radio Astronomy Ob-

servatory, report: “The third x-ray outburst in GRO J1655–40, observed with the Compton Gamma Ray Observatory BATSE (IAUC 6106), was accompanied by a radio outburst observed with the Very Large Array. On Nov. 1, one day after the initial reappearance of the source to BATSE (cf. IAUC 6101), it was at the lowest radio flux levels since initial detection in August: 8, 4, and 3.4 mJy at 1.4, 4.9, and 8.4 GHz. By Nov. 7, it increased its flux levels at the same frequencies to 14, 10, and 10 mJy, followed by a more dramatic increase to 50, 45, and 35 mJy on Nov. 11, and to 80, 76, and 72 mJy on Nov. 15. The next observation on Nov. 18 showed that it had dropped to 31, 25, and 23 mJy. This behavior is similar to that of the first two x-ray/radio events, with a slow increase in radio emission during the rise of the x-ray outburst, and a stronger increase in radio emission associated with the onset of the decay of the x-ray event.”

1994 November 22 Daniel W. E. Green